# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 2.1 Bubbles of Nothing
While Witten’s Bubble of Nothing is in many aspects very simple, it serves as a first step from which we can generalize to more involved bubbles of nothing. In general, starting with a $X_{d+n}=\mathbb{R}^{1,d-1}\times\mathcal{C}_{n}$ Minkowski compactification, the associated Euclidean instanton mediating the bubble of nothing will have $\hat{\mathcal{S}}_{d+n}\simeq\mathbb{S}^{d-1}\times_{\rm W}\mathcal{B}_{n+1}$ topology, with $\mathcal{B}_{n+1}$ the bordism from $\mathcal{C}_{n}$ to nothing, i.e., $\partial\mathcal{B}_{n+1}=\mathcal{C}_{n}$. In the language of (2.5), this results in the following $SO(d-1)$ symmetric ansatz GarciaEtxebarria:2020xsr for the Euclidean solution:
@@
As explained earlier in this section, for Witten’s BoN to exist, it was crucial that the compactification manifold $\mathbb{S}^{1}$ had antiperiodic (i.e., SUSY-breaking) boundary conditions. Furthermore, one would have to argue that the process is not kinematically obstructed and has a finite decay rate. This translates to more general BoN decays through the following two conditions GarciaEtxebarria:2020xsr :
@@
  • Topological Condition: This first requirement translates in all compactification manifolds belonging to the trivial cobordism class. When considering $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ to be $\mathsf{g}$-manifolds, with $\mathsf{g}$ some appropiate structure, then the $\mathsf{g}$-bordism is defined such that $\mathcal{B}_{k+1}$ is also a $\mathsf{g}$-manifold, and the associated $\mathsf{g}$-structure restricted to each of $\partial\mathcal{B}_{k+1}$ components corresponds to that of $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$. The notion of cobordism defines an equivalence relation between compact manifolds, and $[\mathcal{M}_{k}]=[\mathcal{N}_{k}]$. We denote the set of cobordism classes by $\Omega_{k}^{\mathsf{g}}$, which can be shown to have a Abelian group structure, such that $[\mathcal{M}]+[\mathcal{N}]=[\mathcal{M}\sqcup\overline{\mathcal{N}}]$. Through the Swampland Cobordism Conjecture this condition is automatically satisfied by requiring $\Omega^{\rm QG}_{\bullet}=0$. As the complete, non-perturbative description of quantum gravity is not yet known, we will approximate this QG structure by simpler $\mathsf{g}$ whose cobordism groups are better understood. In section 3.3 we comment on the necessary defects that might be required to trivialize non-zero cobordism groups, $\Omega_{k}^{\mathsf{g}}\neq 0\to\Omega_{k}^{\mathsf{g}+{\rm defects}}=0$.
@@
    Translating this topological condition to Witten’s BoN, the circle with antiperiodic boundary conditions $\mathbb{S}^{1}_{a}$ belongs to the trivial class of $\Omega_{1}^{\rm Spin}\simeq\mathbb{Z}_{2}$ Anderson1967:SPIN , with the second class given by the periodic circle $\mathbb{S}^{1}_{p}$. As on principle the later is an allowed compactification manifold, from the Cobordism Conjecture point of view, there must exist either UV defects that can change the spin structure McNamara:2019rup ; Garcia-Etxebarria:2015ota , or either the QG structure $\mathsf{g}={\rm Spin}$ must be refined to $\hat{\mathsf{g}}$ such that $\Omega^{\hat{\mathsf{g}}}_{1}=0$, for instance to $\Omega_{1}^{\rm Spin^{c}}=0$, see Wan:2018bns .
@@
  • Dynamical Condition: The Positive Energy Theorem (PET) Witten:1981mf ; Hertog:2003ru ; Hertog:2003xg states that the ADM mass of asymptotically $\mathcal{M}_{d+n}\simeq\mathbb{R}^{1,{d-1}}\times\mathcal{C}_{n}$ (with $\mathcal{C}_{n}$ compact) spacetimes is bounded from below by 0, with the only spacetime saturating the inequality being precisely $\mathbb{R}^{1,{d-1}}\times\mathcal{C}_{n}$. This prevents any bubble nucleation, since any BoN spacetime will have more energy, and as such the decay will be obstructed. PET relays on the assumption that (1) $\mathcal{M}_{d+n}$ admits an asymptotically covariantly constant spinor (always the case when SUSY is preserved) and (2) the Dominant Energy Condition (DEC), i.e., $-g_{BC}T^{AB}k^{C}$ is causal and future-pointing for all $k^{C}$ causal and future-pointing vectors, holds. While (1) is a topological condition which depends on the global structure of spacetime, (2) depends on the matter and higher-derivative ingredients of our EFT. There is a rich literature on how any of these two conditions can be broken, specially (2) Coleman:1977py ; Curiel:2014zba ; GarciaEtxebarria:2020xsr , so that the PET does not apply.
@@
As for the most part of this paper we will be interested in the qualitative properties of our bordisms to nothing, we will assume that the above conditions hold and thus such bordism can be realizable. In general, the difficulty in building the appropriate BoN solution for a given compactification stems from finding the appropriate bordism solution $h^{\mathcal{B}}_{\alpha\beta}$ to the equations of motion, as well as including the appropriate defects to kill the bordism group $\Omega_{n}^{\mathsf{g}}$. While much effort has been done in the past years in finding explicit BoN constructions Witten:1981gj ; Fabinger:2000jd ; Dine:2004uw ; Horowitz:2007pr ; Yang:2009wz ; Blanco-Pillado:2010xww ; Blanco-Pillado:2010vdp ; Brown:2010mf ; Blanco-Pillado:2011fcm ; Brown:2011gt ; deAlwis:2013gka ; Brown:2014rka ; Blanco-Pillado:2016xvf ; Ooguri:2017njy ; Dibitetto:2020csn ; GarciaEtxebarria:2020xsr ; Bomans:2021ara ; Draper:2021ujg ; Draper:2021qtc ; Petri:2022yhy ; Bandos:2023yyo ; Ookouchi:2024tfz ; Heckman:2024zdo (see also Draper:2023ulp ; Blanco-Pillado:2023hog ; Blanco-Pillado:2023aom ; Delgado:2023uqk ; Friedrich:2024aad for related problems and techniques), most of the solutions feature relatively simple topologies along the bordism direction. One can imagine that as the topology of the internal manifold $\mathcal{C}_{n}$ becomes more involved, so will that of the bordism $\mathcal{B}_{n+1}$. In section 3 we will see that indeed we can obtain plenty of information about these simply from the topology of $\mathcal{C}_{n}$.
@@ 2.2 End of the World Branes
In the above section, we have kept a higher-dimensional description. One can wonder now how Witten’s BoN looks from the $d$-dimensional description once the internal coordinate is integrated out. Dimensionally reducing (2.1) down to $d$-dimensions (see Angius:2022aeq for more details) results in the following metric in the (Euclidean) Einstein frame:
@@
The above behaviors are not unique to Witten’s BoN, but rather universal features of the close distance behavior of End of the World branes. As introduced in Angius:2022aeq , general solutions to $d$-dimensional Einstein gravity coupled to a canonically normalized scalar $\phi$ (generalizable to more scalars and non-canonical metrics) and a potential $V(\phi)$, in the presence of a real codimension-1 boundary (the End of the World Brane), feature the following scalings
@@
Both scalings are given in terms of a critical exponent $\delta$ given by
@@
The example above features a smooth bordism into nothing, with the internal cycle quietly shrinking to a point. In Friedrich:2023tid a general analysis of non-smooth bordisms is presented, with the angular defect at the tip of the bordism being related to the tension of the EotW brane, Vilenkin:1984ib ; Garfinkle:1985hr .
The EotW brane description has been extensively studied from the dynamical cobordism point of view Dudas:2000ff ; Dudas:2002dg ; Dudas:2004nd ; Basile:2018irz ; Antonelli:2019nar ; Basile:2021mkd ; Mourad:2021qwf ; Mourad:2021roa ; Buratti:2021yia ; Buratti:2021fiv ; Angius:2022mgh ; Blumenhagen:2022mqw ; Basile:2022ypo ; Blumenhagen:2023abk ; Angius:2023xtu ; Huertas:2023syg ; Friedrich:2023tid ; Angius:2023uqk ; GarciaEtxebarria:2024jfv ; Huertas:2024mvy ; Angius:2024pqk , and while there is not a one-to-one correspondence between the critical exponent $\delta$ and the UV resolution of the EotW brane, some steps have been performed towards a classification Angius:2024zjv . While in the cobordisms to nothing from Witten’s BoN and other instances in the literature (see Angius:2022aeq for some simple examples) the scalar $\phi$ denotes the volume of the internal manifold, which shrinks to nothing, as we will see in section 4, more involved bordisms have several (canonically normalized) moduli going to infinite distance as we radially traverse our bordism towards its “boundary to nothing”. These scalars need not blow up at the same locus, with the EotW brane being “thick” and having some internal structure, precisely corresponding with the different topology changes which we will describe shortly in 3. This is in no way surprising, as both the EFT description breaking can be seen from the scalars traveling to infinite distance in moduli space, or equivalently from the change in the compactification manifold topology which results in a different lower-dimensional EFT.
@@ 3 Bordism Topology and Morse-Bott Inequalities
As an illustrative example, take $\mathbb{S}^{1}$ together with the oriented bordism. As $\Omega_{1}^{SO}=0$ Thom1954QuelquesPG , there is no obstruction from this side, and actually it is straightforward that $\mathbb{S}^{1}=\partial\mathbb{D}^{2}$. However, more generally the circle is the boundary of any genus-$g$ Riemann surface $\Sigma_{g}$ with a disk removed, $\mathbb{S}^{1}=\partial(\Sigma_{g}-\mathbb{D}^{2})$, with $\mathbb{D}^{2}=\mathbb{S}^{2}-\mathbb{D}^{2}$, see Figure 3. More in general, given some compact $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}$, one can always define $\mathcal{B}_{n+1}^{\prime}=\mathcal{B}_{n+1}\#\mathcal{Y}_{n+1}$, with some appropriate $\mathcal{Y}_{n+1}$ not homeomorphic to $\mathbb{S}^{n+1}$ and to which the structure $\mathsf{g}$ also extends. This way $\mathcal{B}_{n+1}^{\prime}$ and $\mathcal{B}_{n+1}$ will be topologically distinct while both having $\mathcal{C}_{n}$ as boundary. This means that (at least only from topological data), both $\mathbb{S}^{d-1}\times_{\rm W}\mathcal{B}_{n+1}$ and $\mathbb{S}^{d-1}\times_{\rm W}\mathcal{B}_{n+1}^{\prime}$ are two different topologies for BoN decay channels of $\mathbb{R}^{d}\times\mathcal{C}_{n}$, both looking the same far from the bubble wall/EotW brane.
@@
in the local compactification manifold, given by the section or fiber of $\mathcal{B}_{n+1}$. The local theory undergoes in the compact manifold, with different internal cycles opening and closing, until we reach the boundary to nothing. As an illustration, return to the bordisms of the circle $\mathbb{S}^{1}=\partial\mathbb{D}^{2}=\partial(\mathbb{T}^{2}-\mathbb{D}^{2})$, depicted in Figure 3. While in the former case the minimal topology change is $\mathbb{S}^{1}\to\emptyset$ as the circle shrinks to nothing (Witten’s BoN), for the later one undergoes $\mathbb{S}^{1}\to\mathbb{S}^{1}\sqcup\mathbb{S}^{1}\to\mathbb{S}^{1}\to\emptyset$, which involves two extra topology changes.
@@ 3.1 Bounding the topology of the bordism
The first of the two above tasks is the most straightforward, as it will only involve some careful manipulation of long exact sequences. For simplicity, we will restrict to bordisms admitting a continuous deformation to a smooth topological manifold [^6] For example, we will allow for conical singularities, as these could be smoothed-out to a finite curvature region. Some other bordism present in the literature, such as $(\mathbb{Z}/\mathbb{Z}_{2})\times\mathbb{S}_{p}^{1}$ or (×$(\mathbb{R}\times\mathbb{S}_{p}^{1})/\mathbb{Z}_{2}$ for type IIA or IIB string theory on a periodic circle $\mathbb{S}^{1}_{p}$, see McNamara:2019rup , are not. Other bordisms in the same reference include M-theory on non-orientable manifolds. A description of these non-smooth(able) or non-orientable bordisms is out of the reach and assumptions needed to apply the algebraic topology and Morse theory tools we use, so they will not be considered.. Consider for this a $(n+1)$-dimensional smooth manifold $\mathcal{B}_{n+1}$ with boundary $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}\neq\emptyset$. For further simplicity we will assume $\mathcal{B}_{n+1}$ to be orientable.
We start applying by applying the snake lemma, which results in the following long exact sequence:
$\cdots\to H_{k}(\mathcal{C}_{n};\mathbb{Z})\to H_{k}(\mathcal{B}_{n+1};\mathbb{Z})\to H_{k}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\to H_{k-1}(\mathcal{C}_{n};\mathbb{Z})\to\cdots\;,$ (3.2)
where $H_{k}(\mathcal{X};\mathbb{Z})$ is the $k$-th homology group with integer coefficients and $H_{k}(\mathcal{X},\mathcal{Y};\mathbb{Z})$, with $\mathcal{Y}\subseteq\mathcal{X}$, is the relative homology group, whose elements are represented the $n$-chains $\alpha$ of $\mathcal{X}$ whose boundary $\partial\alpha$ is a $(n-1)$-chain of $\mathcal{Y}\subseteq\mathcal{X}$. Applying now Lefschetz duality, $H^{k}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\simeq H_{n+1-k}(\mathcal{B}_{n+1};\mathbb{Z})$, we can rewrite (3.2) as
$\cdots\to H^{n-k}(\mathcal{B}_{n+1};\mathbb{Z})\simeq H_{k+1}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\to H_{k}(\mathcal{C}_{n};\mathbb{Z})\to H_{k}(\mathcal{B}_{n+1};\mathbb{Z})\to\cdots\;.$ (3.3)
@@
note the extra $-1$ in the torsion part. Now, for integral (co)homology, the Universal Coefficient Theorem states that
$H_{k}(\mathcal{M};\mathbb{Z})\simeq\mathbb{Z}^{b_{k}(\mathcal{M})}\otimes T_{k}(\mathcal{M})\;\Longleftrightarrow\;H^{k}(\mathcal{M};\mathbb{Z})\simeq\mathbb{Z}^{b_{k}(\mathcal{M})}\otimes T_{k-1}(\mathcal{M})\;,$ (3.6)
@@
$\cdots\to\mathbb{Z}^{b_{n-k}(\mathcal{B}_{n+1})}\otimes T_{n-k-1}(\mathcal{B}_{n+1})\to\mathbb{Z}^{b_{k}(\mathcal{C}_{n})}\otimes T_{k}(\mathcal{C}_{n})\to\mathbb{Z}^{b_{k}(\mathcal{B}_{n+1})}\otimes T_{k}(\mathcal{B}_{n+1})\to\cdots\;.$ (3.7)
Denoting $t_{k}(\mathcal{M})=\operatorname{rank}T_{k}(\mathcal{M})$, we can conclude
$b_{n-k}(\mathcal{B}_{n+1})+b_{k}(\mathcal{B}_{n+1})$ $\geq b_{k}(\mathcal{C}_{n})=b_{n-k}(\mathcal{C}_{n})$ (3.8a)
$t_{n-k-1}(\mathcal{B}_{n+1})+t_{k}(\mathcal{B}_{n+1})$ $\geq t_{k}(\mathcal{C}_{n})=t_{n-k-1}(\mathcal{C}_{n})\;,$ (3.8b)
for $k=0,\,...,\,n$. As a matter of fact the first inequality is always saturated for the middle Betti numbers when $n$ is even, in the so-called Half lives, half dies Theorem Putman:2022 ,

### v2

@@ 2.1 Bubbles of Nothing
While Witten’s Bubble of Nothing is in many aspects very simple, it serves as a first step from which we can generalize to more involved bubbles of nothing. In general, starting with a $X_{d+n}=\mathbb{R}^{1,d-1}\times\mathcal{C}_{n}$ Minkowski compactification, the associated Euclidean instanton mediating the bubble of nothing will have $\hat{\mathcal{S}}_{d+n}\simeq\mathbb{S}^{d-1}\times_{\rm W}\mathcal{B}_{n+1}$ topology, with $\mathcal{B}_{n+1}$ the bordism from $\mathcal{C}_{n}$ to nothing, i.e., $\partial\mathcal{B}_{n+1}=\mathcal{C}_{n}$. In the language of (2.5), this results in the following SO(d1)SO(d-1) symmetric ansatz García Etxebarria et al. (2020) for the Euclidean solution:
@@
As explained earlier in this section, for Witten’s BoN to exist, it was crucial that the compactification manifold $\mathbb{S}^{1}$ had antiperiodic (i.e., SUSY-breaking) boundary conditions. Furthermore, one would have to argue that the process is not kinematically obstructed and has a finite decay rate. This translates to more general BoN decays through the following two conditions García Etxebarria et al. (2020):
@@
  • Topological Condition: This first requirement translates in all compactification manifolds belonging to the trivial cobordism class. When considering $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$ to be $\mathsf{g}$-manifolds, with $\mathsf{g}$ some appropiate structure, then the $\mathsf{g}$-bordism is defined such that $\mathcal{B}_{k+1}$ is also a $\mathsf{g}$-manifold, and the associated $\mathsf{g}$-structure restricted to each of $\partial\mathcal{B}_{k+1}$ components corresponds to that of $\mathcal{M}_{k}$ and $\mathcal{N}_{k}$. The notion of cobordism defines an equivalence relation between compact manifolds, and $\mathcal{M}_{k}]=[\mathcal{N}_{k}]$. We denote the set of cobordism classes by $\Omega_{k}^{\mathsf{g}}$, which can be shown to have a Abelian group structure, such that $\mathcal{M}]+[\mathcal{N}]=[\mathcal{M}\sqcup\overline{\mathcal{N}}]$. Through the Swampland Cobordism Conjecture this condition is automatically satisfied by requiring $\Omega^{\rm QG}_{\bullet}=0$. As the complete, non-perturbative description of quantum gravity is not yet known, we will approximate this QG structure by simpler $\mathsf{g}$ whose cobordism groups are better understood. In section 3.3 we comment on the necessary defects that might be required to trivialize non-zero cobordism groups, $\Omega_{k}^{\mathsf{g}}\neq 0\to\Omega_{k}^{\mathsf{g}+{\rm defects}}=0$.
@@
    Translating this topological condition to Witten’s BoN, the circle with antiperiodic boundary conditions $\mathbb{S}^{1}_{a}$ belongs to the trivial class of $\Omega_{1}^{\rm Spin}\simeq\mathbb{Z}_{2}$ Anderson et al. (1967), with the second class given by the periodic circle $\mathbb{S}^{1}_{p}$. As in principle the later is an allowed compactification manifold, from the Cobordism Conjecture point of view, there must exist either UV defects that can change the spin structure McNamara and Vafa (2019); García-Etxebarria et al. (2015), or either the QG structure $\mathsf{g}={\rm Spin}$ must be refined to $\hat{\mathsf{g}}$ such that $\Omega^{\hat{\mathsf{g}}}_{1}=0$, for instance to $\Omega_{1}^{\rm Spin^{c}}=0$, see Wan and Wang (2019).
@@
  • Dynamical Condition: The Positive Energy Theorem (PET) Witten (1981); Hertog et al. (2003, 2004) states that the ADM mass of asymptotically $\mathcal{M}_{d+n}\simeq\mathbb{R}^{1,{d-1}}\times\mathcal{C}_{n}$ (with $\mathcal{C}_{n}$ compact) spacetimes is bounded from below by 0, with the only spacetime saturating the inequality being precisely $\mathbb{R}^{1,{d-1}}\times\mathcal{C}_{n}$. This prevents any bubble nucleation, since any BoN spacetime will have more energy, and as such the decay will be obstructed. PET relays on the assumption that (1) $\mathcal{M}_{d+n}$ admits an asymptotically covariantly constant spinor (always the case when SUSY is preserved) and (2) the Dominant Energy Condition (DEC), i.e., $g_{BC}T^{AB}k^{C}$ is causal and future-pointing for all $k^{C}$ causal and future-pointing vectors, holds. While (1) is a topological condition which depends on the global structure of spacetime, (2) depends on the matter and higher-derivative ingredients of our EFT. There is a rich literature on how any of these two conditions can be broken, specially (2) Coleman (1977); Curiel (2017); García Etxebarria et al. (2020), so that the PET does not apply.
@@
As for the most part of this paper we will be interested in the qualitative properties of our bordisms to nothing, we will assume that the above conditions hold and thus such bordism can be realizable. In general, the difficulty in building the appropriate BoN solution for a given compactification stems from finding the appropriate bordism solution $h^{\mathcal{B}}_{\alpha\beta}$ to the equations of motion, as well as including the appropriate defects to kill the bordism group $\Omega_{n}^{\mathsf{g}}$. While much effort has been done in the past years in finding explicit BoN constructions Witten (1982a); Fabinger and Horava (2000); Dine et al. (2004); Horowitz et al. (2008); Yang (2010); Blanco-Pillado and Shlaer (2010); Blanco-Pillado et al. (2010); Brown and Dahlen (2011); Blanco-Pillado et al. (2012); Brown and Dahlen (2012); de Alwis et al. (2013); Brown (2014); Blanco-Pillado et al. (2016); Ooguri and Spodyneiko (2017); Dibitetto et al. (2020); García Etxebarria et al. (2020); Bomans et al. (2022); Draper et al. (2021b, a); Petri (2022); Bandos et al. (2023); Ookouchi et al. (2024); Heckman and Hübner (2024) (see also Draper et al. (2023); Blanco-Pillado et al. (2024b, a); Delgado (2024); Friedrich and Hebecker (2024) for related problems and techniques), most of the solutions feature relatively simple topologies along the bordism direction. One can imagine that as the topology of the internal manifold $\mathcal{C}_{n}$ becomes more involved, so will that of the bordism $\mathcal{B}_{n+1}$. In section 3 we will see that indeed we can obtain plenty of information about these simply from the topology of $\mathcal{C}_{n}$.
@@ 2.2 End of the World Branes
In the above section, we have kept a higher-dimensional description. One can wonder now how Witten’s BoN looks from the $d$-dimensional description once the internal coordinate is integrated out. Dimensionally reducing (2.1) down to $d$-dimensions (see Angius et al. (2022a) for more details) results in the following metric in the (Euclidean) Einstein frame:
@@
The above behaviors are not unique to Witten’s BoN, but rather universal features of the close distance behavior of End of the World branes. As introduced in Angius et al. (2022a), general solutions to $d$-dimensional Einstein gravity coupled to a canonically normalized scalar $\phi$ (generalizable to more scalars and non-canonical metrics) and a potential $V(\phi)$, in the presence of a real codimension-1 boundary (the End of the World Brane), feature the following scalings
@@
Both scalings are expressed in terms of a critical exponent $\delta$ given by
@@
The example above features a smooth bordism into nothing, with the internal cycle quietly shrinking to a point. In Friedrich et al. (2024) a general analysis of non-smooth bordisms is presented, with the angular defect at the tip of the bordism being related to the tension of the EotW brane, Vilenkin (1985); Garfinkle (1985).
The EotW brane description has been extensively studied from the dynamical cobordism point of view Dudas and Mourad (2000); Dudas et al. (2003, 2005); Basile et al. (2019); Antonelli and Basile (2019); Basile (2021); Mourad and Sagnotti (2021a, b); Buratti et al. (2021b, a); Angius et al. (2022b); Blumenhagen et al. (2022); Basile et al. (2022); Blumenhagen et al. (2023); Angius et al. (2023); Huertas and Uranga (2023); Friedrich et al. (2024); Angius et al. (2024a); García Etxebarria et al. (2024); Huertas and Uranga (2024); Angius et al. (2024b), and while there is not a one-to-one correspondence between the critical exponent $\delta$ and the UV resolution of the EotW brane, some steps have been performed towards a classification Angius (2024). While in the cobordisms to nothing from Witten’s BoN and other instances in the literature (see Angius et al. (2022a) for some simple examples) the scalar $\phi$ denotes the volume of the internal manifold, which shrinks to nothing, as we will see in section 4, more involved bordisms have several (canonically normalized) moduli going to infinite distance as we radially traverse our bordism towards its “boundary to nothing”. These scalars need not blow up at the same locus, with the EotW brane being “thick” and having some internal structure, precisely corresponding with the different topology changes which we will describe shortly in 3. This is in no way surprising, as both the EFT description breaking can be seen from the scalars traveling to infinite distance in moduli space, or equivalently from the change in the compactification manifold topology which results in a different lower-dimensional EFT.
@@ 3 Bordism Topology and Morse-Bott Inequalities
As an illustrative example, take $\mathbb{S}^{1}$ together with the oriented bordism. As $\Omega_{1}^{SO}=0$ Thom (1954), there is no obstruction from this side, and actually it is straightforward that $\mathbb{S}^{1}=\partial\mathbb{D}^{2}$. However, more generally the circle is the boundary of any genus-$g$ Riemann surface $\Sigma_{g}$ with a disk removed, $\mathbb{S}^{1}=\partial(\Sigma_{g}-\mathbb{D}^{2})$, with $\mathbb{D}^{2}=\mathbb{S}^{2}-\mathbb{D}^{2}$, see Figure 3. More in general, given some compact $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}$, one can always define $\mathcal{B}_{n+1}^{\prime}=\mathcal{B}_{n+1}\#\mathcal{Y}_{n+1}$, with some appropriate $\mathcal{Y}_{n+1}$ not homeomorphic to $\mathbb{S}^{n+1}$ and to which the structure $\mathsf{g}$ also extends. This way $\mathcal{B}_{n+1}^{\prime}$ and $\mathcal{B}_{n+1}$ will be topologically distinct while both having $\mathcal{C}_{n}$ as boundary. This means that (at least only from topological data), both $\mathbb{S}^{d-1}\times_{\rm W}\mathcal{B}_{n+1}$ and $\mathbb{S}^{d-1}\times_{\rm W}\mathcal{B}_{n+1}^{\prime}$ are two different topologies for BoN decay channels of $\mathbb{R}^{d}\times\mathcal{C}_{n}$, both looking the same far from the bubble wall/EotW brane.
@@
in the local compactification manifold, given by the section or fiber of $\mathcal{B}_{n+1}$. The local theory undergoes a series of changes in the compact manifold, with different internal cycles opening and closing, until we reach the boundary to nothing. As an illustration, return to the bordisms of the circle $\mathbb{S}^{1}=\partial\mathbb{D}^{2}=\partial(\mathbb{T}^{2}-\mathbb{D}^{2})$, depicted in Figure 3. While in the former case the minimal topology change is $\mathbb{S}^{1}\to\emptyset$ as the circle shrinks to nothing (Witten’s BoN), for the later one undergoes $\mathbb{S}^{1}\to\mathbb{S}^{1}\sqcup\mathbb{S}^{1}\to\mathbb{S}^{1}\to\emptyset$, which involves two extra topology changes.
@@ 3.1 Bounding the topology of the bordism
The first of the two above tasks is the most straightforward, as it will only involve some careful manipulation of long exact sequences. For simplicity, we will restrict to bordisms admitting a continuous deformation to a smooth topological manifold [^6] For example, we will allow for conical singularities, as these could be smoothed-out to a finite curvature region. Some other bordism present in the literature, such as $(\mathbb{Z}/\mathbb{Z}_{2})\times\mathbb{S}_{p}^{1}$ or $(\mathbb{R}\times\mathbb{S}_{p}^{1})/\mathbb{Z}_{2}$ for type IIA or IIB string theory on a periodic circle $\mathbb{S}^{1}_{p}$, see McNamara and Vafa (2019), are not. Other bordisms in the same reference include M-theory on non-orientable manifolds. A description of these non-smooth(able) or non-orientable bordisms is out of the reach and assumptions needed to apply the algebraic topology and Morse theory tools we use, so they will not be considered.. Consider for this a $(n+1)$-dimensional smooth manifold $\mathcal{B}_{n+1}$ with boundary $\mathcal{C}_{n}=\partial\mathcal{B}_{n+1}\neq\emptyset$. For further simplicity we will assume $\mathcal{B}_{n+1}$ to be orientable.
We start by applying the zig-zag lemma, which results in the following long exact sequence:
$\cdots\xrightarrow{\partial}H_{k}(\mathcal{C}_{n};\mathbb{Z})\xrightarrow{i_{\ast}}H_{k}(\mathcal{B}_{n+1};\mathbb{Z})\xrightarrow{j_{\ast}}H_{k}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\xrightarrow{\partial}H_{k-1}(\mathcal{C}_{n};\mathbb{Z})\xrightarrow{i_{\ast}}\cdots\;,$ (3.2)
where $H_{k}(\mathcal{X};\mathbb{Z})$ is the $k$-th homology group with integer coefficients and $H_{k}(\mathcal{X},\mathcal{Y};\mathbb{Z})$, with $\mathcal{Y}\subseteq\mathcal{X}$, is the relative homology group, whose elements are represented by the $n$-chains $\alpha$ of $\mathcal{X}$ whose boundary $\partial\alpha$ is a (n1)(n-1)-chain of $\mathcal{Y}\subseteq\mathcal{X}$. Applying now Lefschetz duality, [^7] In the pairings between homology and cohomology groups through Lefschetz duality, cohomology is to be taken with compact support, $H_{k}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\simeq H^{n+1-k}_{\mathbf{c}}(\mathcal{B}_{n+1};\mathbb{Z})$ and $H^{k}_{\mathbf{c}}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\simeq H_{n+1-k}(\mathcal{B}_{n+1};\mathbb{Z})$, Maunder (1996). However, since $\mathcal{B}_{n+1}$ is compact, we do not need to worry about this subtlety, and can consider de Rham cohomology. $H_{k}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\simeq H^{n+1-k}(\mathcal{B}_{n+1};\mathbb{Z})$, we can rewrite (3.2) as
$\cdots\xrightarrow{j_{\ast}}H^{n-k}(\mathcal{B}_{n+1};\mathbb{Z})\simeq H_{k+1}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})\xrightarrow{\partial}H_{k}(\mathcal{C}_{n};\mathbb{Z})\xrightarrow{i_{\ast}}H_{k}(\mathcal{B}_{n+1};\mathbb{Z})\xrightarrow{j_{\ast}}\cdots\;.$ (3.3)
@@
note the extra 1-1 in the torsion part. Now, for integral (co)homology, the Universal Coefficient Theorem states that
$H_{k}(\mathcal{M};\mathbb{Z})\simeq\mathbb{Z}^{b_{k}(\mathcal{M})}\oplus T_{k}(\mathcal{M})\;\Longleftrightarrow\;H^{k}(\mathcal{M};\mathbb{Z})\simeq\mathbb{Z}^{b_{k}(\mathcal{M})}\oplus T_{k-1}(\mathcal{M})\;,$ (3.6)
@@
$\cdots\to\mathbb{Z}^{b_{n-k}(\mathcal{B}_{n+1})}\oplus T_{n-k-1}(\mathcal{B}_{n+1})\to\mathbb{Z}^{b_{k}(\mathcal{C}_{n})}\oplus T_{k}(\mathcal{C}_{n})\to\mathbb{Z}^{b_{k}(\mathcal{B}_{n+1})}\oplus T_{k}(\mathcal{B}_{n+1})\to\cdots\;.$ (3.7)
Denoting $t_{k}(\mathcal{M})={\rm rank}\,T_{k}(\mathcal{M})$, we can conclude [^8] We can study the free and torsional ranks separately, as there are no torsional $k$-cycles acting as boundary of torsion-free cycles in $H_{k+1}(\mathcal{B}_{n+1},\mathcal{C}_{n};\mathbb{Z})$. To see this, we tensor the (3.2) long exact sequence by $\mathbb{Q}$, which acts as an exact functor, getting rid of the torsion parts:
[…]
$b_{k}(\mathcal{B}_{n+1})+b_{n-k}(\mathcal{B}_{n+1})$ $\geq b_{k}(\mathcal{C}_{n})=b_{n-k}(\mathcal{C}_{n})$ (3.8a)
$t_{k}(\mathcal{B}_{n+1})+t_{n-k-1}(\mathcal{B}_{n+1})$ $\geq t_{k}(\mathcal{C}_{n})=t_{n-k-1}(\mathcal{C}_{n})\;,$ (3.8b)
for k=0,,nk=0,\,...,\,n. As a matter of fact the first inequality is always saturated for the middle Betti numbers when $n$ is even, in the so-called Half lives, half dies Theorem Putman (2022),